OpenStudy (anonymous):

What is the simplified form of 15x^6/20y^5*6y^2/5x^4?

OpenStudy (anonymous):

\[15x ^{6}/20y ^{5}\times6y ^{2}/5x ^{4}\]

OpenStudy (pfenn1):

\[\frac{15x^6}{20y^5}*\frac{6y^2}{5x^4}\]Can you simplify \[\frac{15x^6}{5x^4}*\frac{6y^2}{20y^5}\]

OpenStudy (anonymous):

Idk how to :/

OpenStudy (anonymous):

I don't understand this at all

OpenStudy (anonymous):

I'm confused.

OpenStudy (pfenn1):

Ok. All I was trying to do was to rearrange the terms in the expression in a way as to make them easier to recognize. However, I have confused you, for which I must apologize. Let's go back to the original expression\[\frac{15x^6}{20y^5}*\frac{6y^2}{5x^4}\]What do you get when you multiply out this expression?